A Bibliography of Publications in ACM SIGGRAPH Computer Graphics: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 585 1640, +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
                    WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.09

Title word cross-reference

+ [CH11]. 3 [Her10].

academia [Cha10]. advanced [Hsi10]. age [Wes10c]. algorithm [DD10]. Animation [Fer11, Van10, Cha10]. ANIMIAMI [Fer11]. anisotropic [KSKK10]. approach [Ma10, Van10]. Art [CH11, Wes10b]. assess [Wes10b]. assessment [Wes10c].

based [Sch11, Wes10b]. BRDF [KSKK10].

Carlo [KSKK10]. collaboration [Wes11]. computer [Her10, Wes10c, Wes11]. Conference [Fer11]. content [Her10]. Coordination [Wes11]. creating [Her10]. creativity [CH11]. Cross [Dav10].

Cross-media [Dav10]. culture [DD10].

D [Her10]. data [Ma10].

Early [Kee10]. earthquakes [Hsi10]. Ending [Her10]. evaluation [Wes10c].

Festival [Fer11]. fitting [KSKK10]. flame [DD10]. flowers [Wes10a]. frontier [CH11].

graphics [Her10, Wes10c, Wes11].

Haley [Hal10].

impact [Wes11, Wri11]. industry [Cha10]. innovation [CH11]. interactive [Sch11]. Internet [Wes10b].

large [Ma10]. lunar [Wri11]. media [Dav10]. Meet [Hal10]. memories [Kee10]. MK [Hal10]. modal [Sch11]. model [KSKK10]. Monte [KSKK10]. multi [Sch11]. multi-modal [Sch11].

open [DD10].

perceptions [Kee10]. Persistence [Wes10a]. physically [Sch11]. Preparing [Wri11]. process [Wes10b]. production [Her10].

quality [Her10].

real [Her10]. real-time [Her10]. remote [Ma10]. rendering [Her10, KSKK10]. renegade [Van10]. research [Cha10, Wes10c].

skills [Wes10b]. software [Her10]. sound [Sch11]. source [DD10]. stepping [Her10]. stories [Dav10]. student [Wes10b].

technology [CH11, Van10]. time [Her10]. tradeoff [Her10].

Understanding [Hsi10]. Using [Wes10b].

Visual [Wes10c]. visualization [Hsi10, Ma10]. volume [Ma10]. vs [Her10].

wild [Wes10a]. workflow [Her10]. workspace [Sch11].

References

Chestney-Harvey:2011:ATN

[CH11] Kim Chestney-Harvey. Art + technology: the new frontier of creativity and innovation. Computer Graphics, 45(1):2:1–2:??, February 2011. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Chavez:2010:AIA

[Cha10] Mark Chavez. From the animation industry to academia and research. Computer Graphics, 44(1): 5:1–5:??, February 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Davidson:2010:CMS

[Dav10] Drew Davidson. Cross-media stories. Computer Graphics, 44(3): 7:1–7:??, August 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Draves:2010:FAO

[DD10] Scott Draves and Isabel Walcott Draves. The flame algorithm and its open source culture. Computer Graphics, 44(3): 3:1–3:??, August 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ferrazza:2011:AAF

[Fer11] Mauricio Ferrazza. ANIMIAMI animation festival and conference. Computer Graphics, 45(1):5:1– 5:??, February 2011. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Haley:2010:MMH

[Hal10] MK Haley. Meet MK Haley. Computer Graphics, 44(3): 6:1–6:??, August 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Herrera:2010:ETT

[Her10] Alex Herrera. Ending the tradeoff of time vs. quality when creating 3D computer graphics content: stepping up the production workflow with real-time rendering software. Computer Graphics, 44(1): 2:1–2:??, February 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hsieh:2010:UEA

[Hsi10] Tung-Ju Hsieh. Understanding earthquakes with advanced visualization. Computer Graphics, 44(1): 4:1–4:??, February 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Keever:2010:EMN

[Kee10] Kim Keever. Early memories & new perceptions. Computer Graphics, 44(3):2:1–2:??, August 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Kurt:2010:ABM

[KSKK10] Murat Kurt, László Szirmay-Kalos, and Jaroslav Křivánek. An anisotropic BRDF model for fitting and Monte Carlo rendering. Computer Graphics, 44(1): 3:1–3:??, February 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ma:2010:NAR

[Ma10] Kwan-Liu Ma. A new approach to remote visualization of large volume data. Computer Graphics, 44 (3):5:1–5:??, August 2010. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Schroeder:2011:IMM

[Sch11] Benjamin Schroeder. An interactive, multi-modal workspace for physically based sound. Computer Graphics, 45(1):3:1–3:??, February 2011. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

VanCitters:2010:ATR

[Van10] Darrell Van Citters. Animation and technology: a renegade approach. Computer Graphics, 44(3): 4:1–4:??, August 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2010:PWF

[Wes10a] Hans Westman. Persistence of the wild flowers. Computer Graphics, 44(3):1:1–1:??, August 2010. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2010:UIB

[Wes10b] Hans Westman. Using an Internet based process to assess student art skills. Computer Graphics, 44(1): 6:1–6:??, February 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2010:VRE

[Wes10c] Hans Westman. Visual research, evaluation and assessment in the age of computer graphics. Computer Graphics, 44(1):1:1–1:??, February 2010. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2011:CCI

[Wes11] Hans Westman. Coordination, collaboration and the impact of computer graphics. Computer Graphics, 45(1):1:1–1:??, February 2011. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wright:2011:PLI

[Wri11] Ernie Wright. Preparing for a lunar impact. Computer Graphics, 45(1):4:1–4:??, February 2011. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.